OpenStudy (anonymous):

can someone please explain the difference between sine, cosine and tangent. and explain how i can find them. also give examples.

OpenStudy (anonymous):

http://www.mathwarehouse.com/trigonometry/sine-cosine-tangent.php This is a really good website.

OpenStudy (elonasushchik):

Well in geometry you would use the trig functions to find out triangle sides or angles Say you have a right triangle and your using trig functions say that you are looking for the base and they give you a side. You Use SIN on an angle. If the angle is 42 you would do SIN 42 = O/H which means the side OPPOSITE of the Angle 42 and the Hypotenuse. You use COS 42 = A/H which means the side length closes to 42 and the Hypotenuse You use TAN 42 = O/A which means opposite side length of the angle and the Adjacent side length of the angle.

OpenStudy (elonasushchik):

@lostsilverfangirl Medal and fan tx?!

OpenStudy (anonymous):

http://www.rasmus.is/uk/t/F/Su30k3.htm And here's another website too.

OpenStudy (anonymous):

thank u . both of u

OpenStudy (anonymous):

You're Welcome:)